OpenStudy (anonymous):

Which of the following is the general solution of the differential equation dy dx equals the quotient of 8 times x and y?

OpenStudy (anonymous):

https://gyazo.com/16816aebcb80f28e1fb069e65da2251f

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (jango_in_dtown):

kindly type the question in the comment section in mathematical form

OpenStudy (jango_in_dtown):

\[\frac{ dy }{ dx}=\frac{ 8x }{ y} \]

OpenStudy (jango_in_dtown):

implies \[ydy=8xdx\]

OpenStudy (jango_in_dtown):

integrating we get \[\frac{ y^2 }{ 2 }=8\frac{ x^2 }{ 2 }+\frac{ C }{ 2 }\]

OpenStudy (jango_in_dtown):

implies \[y^2=8x^2+C\]

OpenStudy (jango_in_dtown):

@Ephemera